OpenStudy (anonymous):

A line passes through the points (1, 2) and (4, 4). What is the slope of the line?

OpenStudy (anonymous):

-3/-2 ?

OpenStudy (haseeb96):

M=y2-y1/x2-x1 Here (x1,y1)=(1,2) (X2,y2)=(4,4) Just put them into the formula and get the answer

OpenStudy (anonymous):

\[\frac{4-2}{4-1}=\frac{2}{3} \]

OpenStudy (triciaal):

the slope of a line is the ratio of the change in y with the change in x to find the change get the difference in the values given 2 points (x1, y1) and (x2, y2) the slope is (y2 - y1)/ (x2-x1)